OpenStudy (anonymous):

18+27divided by 3^2= 21 can anyone explain how to calculate this without a calculator

OpenStudy (anonymous):

ik the answer is 21 i got it using the calculator but how would i do this on paper ?

OpenStudy (anonymous):

18+27=45. 3^2=9. 45/9=5. the answer should be 5.

OpenStudy (anonymous):

i got 21 on my scientific calculator 18+27/3^2

OpenStudy (anonymous):

Oh, ok. I see what you're doing 27/9=3+18=21

OpenStudy (anonymous):

\[18+21/3^{2}\]\[18+21/9\]\[18+3\]

OpenStudy (anonymous):

was that the correct way of solving

OpenStudy (anonymous):

Yeah. I thought it was \[\frac{ 18+27 }{ 3^{2} }\]but that's not the equation. So yes, 21 is the right answer.

OpenStudy (anonymous):

ok can you correct me on this

OpenStudy (anonymous):

1over 2 - 7 over 8 = - 3 over 8

OpenStudy (anonymous):

srry I'm on a mac dnt know how to put fractions like you

OpenStudy (anonymous):

You're right. 21 is the right answer. 3^2=9. 27/9=3. Then 18+3=21

OpenStudy (anonymous):

It depends if your question is \[\frac{ 18+27 }{ 3^{2} }~~~or~~~18+\frac{ 27 }{ 3^{2} }\]

OpenStudy (ann1987):

The problem should be written like this: 18 + ( 27/9) should equal 21 18 + 3 should equal 21 21 does equal 21

OpenStudy (anonymous):

no the equation is 18 +27 (division sign) 3^2

OpenStudy (anonymous):

and for this equation did i do it correct? 1/2 - 7/8

OpenStudy (anonymous):

i got -3/8